OpenStudy (anonymous):

[8.02] Determine the factors of 15x2 + 3xy + 10x + 2y. (3x + 2)(5x + y) (5x + y)(2x + 3) (3x + y)(5x + 2) (5x + 3)(2x + y)

OpenStudy (anonymous):

Help me please and thank you

OpenStudy (ybarrap):

Let's take the common factors from \(15x^2\) and \(3xy\) first: $$ 15x^2+3xy=3x(5x^2+y) $$ Do you see that?

OpenStudy (ybarrap):

Next, let's find the common factors of 10x and 2y. Do you see it?

OpenStudy (anonymous):

yes

OpenStudy (ybarrap):

tell me

OpenStudy (anonymous):

i see it but how do i solve it

OpenStudy (ybarrap):

Correction to the above: $$ 15x^2+3xy=3x(5x+y) $$ For the next two terms, we have: $$ 10x+2y=2(5x+y) $$ Put it all together then we get $$ 3x(5x+y)+2(5x+y) $$ Do you see the next step?

OpenStudy (anonymous):

yea

OpenStudy (ybarrap):

We have something in common with the \(3x \) and the 2: It's the \( 5x+y\) term. So bring out the \( 5x+y\) term and we have: $$ (5x+y)(3x+2) $$ That it!

OpenStudy (anonymous):

oh ok thank you

OpenStudy (ybarrap):

If you can find like terms between any two pairs, try it and then see if the next two terms have the same factors in common. If that doesn't work, there are other things to try.

OpenStudy (ybarrap):

You're welcome and Good luck!